OpenStudy (abmon98):

A stone is dropped into the water from the roof of a building 44 m above the water. Another stone is thrown vertically down 1.0 s after the first is dropped. Both stones strike the water at the same time. What was the initial speed of the second stone?

OpenStudy (abmon98):

Here is my working I set the displacement to be -44 m because its below the projection level, a=gm/s^2 and the initial velocity to be 0 m/s then i used v^2=u^2+2as v=-29.664 t=2.966 if the other stone is dropped 1 second after the 1st stone then the total time it takes to strike the water is 2.966-1=1.966 a=g, do we make an assumption that the second stone was dropeed from a height of 44 m?